please help me in this problem set! An organization wants to sell tickets to a concert. it plans on selling 300 reserved-seat tickets and 150 tickets at the door. the price of a reserved-seat ticket is to be $2 more than a ticket at the door. if the organization wants to collect at least $3,750, what is the maximum it can charge for a reserved-seat ticket?
Answer by mananth(16949)

300 reserved-seat tickets
150 tickets at the door.
..
ticket at door be x
price of a reserved-seat ticket is to be $2 more = x+2
collection $3,750,
...
150x+300(x+2)=3750
150x+300x+600=3750
450x=3750-600
450x=3150
x=$ 70 at the door price
reserved = $ 72